UT 20140118 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards
Telescope Operator: Sallenson
LBTO Support: TSargent
Visitors: CVeillet, DSummers

Plan for the night

- LS-alignment of two stars
- Ima movement due to flexure
- Long spectroscopical setup for efficiency

Summary

- GEIRS w/o problems entire night

- SW stable (only telescope service hickupped only 3 times)

- Lost 2hrs due to collimation issues (at excellent seeing)

- LS-acq. tests, CofR -determination

- Throughput determination at detector center

- Deep spectroscopical setup

Preparations

- Field aligned dx,y=-1300,4500
- Dave took calibrations this morning
- Alexander put in a new SW-version, Ingo a modifed RTD

Details

All times are in UT, files are luci2.20140117.iiii

Files #1-149 are test data taken by Alexander

Flatfields

00:41
H2: #150-179, 6sec, 16-8k
FeII: #180-209 , 5sec, 18-3k
J_high: #210-239, 6sec, 20-6k

00:40 Sunset, weather photometric, low wind, hum = 10%, T=2deg

01:10 Steve checks pointing & collimates @ BS9107, seeing is fantastic, 0.3" on guider

01:28 BS9107, LS 1.5", script not accepted, restart GUI, OK, alignment 2stars on slit
using now proper CofR as input, sky #240, scr #241, ls #242, alignment,
verification #243 2 stars on slit but the wrong ones, only translation, no rotation
although rotation was properly transported to the telescope, on guider GS first moved down
indicating rotation, catched, after that GS moved up and was catched again
should in principle GS should alwas be directly at hotspot

Redid the experiment, #244-247, this time there was only one jump of GS at guider,
offset & translation was properly applied, not clear what happened, roll back to
previous SW-version in order to check if new telescope service was the reason

Steve cleared mirror, collimation went into bananas, restart GCS

Redid the experiment #248-254, same behaviour at guider, roll back to latest SW-version

03:05 BS9107, PA=225, recheck LS-acq. of 2 stars, sky #255, src #256, LS #256,
aligned, verification #257, no rotation?

03:26 redo, this time with bigger rotation, rotation/no rotation, #258-267
YAK, there is a PA=225 for slit verification in slit, of course, can't work

04.01 BS9107, modified script, #268-271, rotation OK, offset in x and y by 14pix???

04:15 TNO, J #272-274. Heavy elongation on images. Correcting collimation and pointing on a bright star and restarting

04:30 TNO, J 272-281. Collimation is running away again. Clear there is a secondary problem. Calling Doug Miller....

05:00 TNO, #282-290, seeing is exiting, image quality on J-band images a desaster, the appearance of the PSF changes
dramatically from frame to frame

05:16 Decided to stop after set of J-band images, Steve power-cycles secondary, cleared active optics

05:23 TNO, J, #291, see if collimation is OK now, does still look bad, continue to see how this developes, Ks, #292-293, stopped

05:35 Restart GCS, TNO, J, #294, problem still present

05:39 Everthing cleared now, TNO, J #295, not better

05:47 TNO again, this time at diffferent PA and GS, J #296-299, somewhat better, stopped here

06:03 Ima movement, M67, AM=1.2, PA=0, #305, Steve has to correct pointing, restart, PA=0, 90, 180, 270 #311-314

06:32 LS-acquisition 2 stars, sky #315, scr #316, LS #317, aligned, verification #318/319, rotation OK,
but stars of by dx=-9 pix, dy = 4pix, CofR wrongly calculated?

06:32 Redo with old values for CofR (1023,1023), sky #320, src #321, ls #322, aligned,
verification #323, dx=+3, dy=+25

07:18 Ima movement, #330 crap, M67, AM=1.1, PA=0, 90, 180, 270 #331-334

07:48 Redo with new values for CofR (1071,1011), sky #340, src #341, LS #342, nada

08:47 Redo with new values for CofR (974/1062), offset still dx=-9pix, dy=+1 pix
Problem seems to be still the CofR (and the tilt of the slit)

09:43 New CoofR (980/1055), last 2*acq. for tonight, sky #361, src #362, LS #363, verification #364, dx=dy=11pix

09:55 Throughput BS9142, sky #367, src #368, LS, #369, verification #370 OK, spectra #371-405, seeing still excellent

11:15 Throughput tellurics, #406-431, had to stop/start telescope service once

11:56 Kurk GDNPHB31, stop/start telescope service again, sky #432 src #433,MOSview #434, aligned,
verification #435, looks good, moving from Ks to K , FW2-error, took a while to recover,
sci started at 12:44, will do a couple just to see how it goes, #436-445

13:42 Telluric for GDNPHB31, yak, Telluric provided is below horizon (at 22hrs)
Tried to go to HIP60758 on the fly and to take spectra.everything manually..needs to be check. #446-459
Seems something is there, N1.8 camea, slit 107 is #454-456, slit 111 is #457-459
Last spectrum taken 5min before sunrise

Too late for flats, but cirrus outside anyway

14:11 Closing

Daytime calibrations

UT 20140118 LUCI2 Com3 Night3 DThompson
luci2.20140117.NNNN.fits (yes, the date is still incorrect)

### Spectroscopic calibrations for mask M007GDNPHB31

N180 GDNPHB31 G210@2.225um clear+K FCon CUin
0460 0464 5*[1x 4.00s LIR normal] Lamps OFF ~150ADU thermal bkgnd
0465 0469 5*[1x 2.75s LIR normal] Lamps OFF anticipating Halo1 brightness
0470 0474 5*[1x 2.75s LIR normal] Halo1 ON 36k ADU
0475 0479 5*[1x 2.75s LIR normal] Halo2 ON 1.3k ADU
0480 0484 5*[1x 4.00s LIR normal] Halo2 ON 2k ADU
0485 0489 5*[1x20.00s LIR normal] Halo2 ON 8.6k ADU
0490 0494 5*[1x20.00s LIR normal] Lamps OFF 670 ADU

0495 0499 5*[1x10.00s LIR normal] Lamps OFF 400 ADU
0500 1*[1x10.00s LIR normal] Argon ON
0501 1*[1x10.00s LIR normal] Neon ON
0502 1*[1x10.00s LIR normal] Xenon ON not many lines in K

### Sieve mask in our substrate material (sieve150mu)

N180 GDNPHB31sieve150mu mirror clear+K FCon CUin
0503 1*[5x 2.75s LIR integr] FW1=blind
0504 1*[5x 2.75s LIR integr] Lamps OFF FW1=clear 7k? ADU in spots
0505 1*[5x 2.75s LIR integr] Halo3 ON 16k? ADU in spots
N375 0506 1*[5x 2.75s LIR integr] Halo3 ON 12k ADU in spots
0507 1*[5x 2.75s LIR integr] Halo3 OFF 7k ADU in spots
0508 1*[5x 2.75s LIR integr] Halo3 OFF FW1=blind

### xtalk mask (dthompson, redundant design)

0509 1*[5x 2.75s LIR integr] Halo3 OFF FW1=blind
0510 1*[5x10.00s LIR integr] Halo3 OFF FW1=blind
0511 1*[5x 2.75s LIR integr] Halo3 OFF FW1=clear
0512 1*[5x10.00s LIR integr] Halo3 OFF FW1=clear

0513 1*[5x 2.75s LIR integr] Halo3 ON FW1=clear
0514 1*[5x10.00s LIR integr] Halo3 ON FW1=clear
0515 1*[3x30.00s LIR integr] Halo3 ON FW1=clear right around saturation
0516 1*[1x100.00s LIR normal] Halo3 ON FW1=clear

0517 0518 1*[5x 2.75s LIR integr] Halo2 ON FW1=clear
0519 1*[5x10.00s LIR integr] Halo2 ON FW1=clear
0520 1*[3x30.00s LIR integr] Halo2 ON FW1=clear

0521 0522 1*[5x 2.75s LIR integr] Halo1 ON FW1=clear
0523 1*[5x10.00s LIR integr] Halo1 ON FW1=clear
0524 1*[3x30.00s LIR integr] Halo1 ON FW1=clear

Reconfigure with blind mask, blind+clear filters, mirror, N375, Calib unit out, FC off

While mechanisms were moving, start:
0525 0534 1*[1x60.00s MER normal] persistence spots several tens of ADU peak after 10 minutes
0535 0554 1*[1x300.00s MER normal]

-- JochenHeidt - 18 Jan 2014
Topic revision: r9 - 20 Jan 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback